# Microscopic origin of the nemato-elastic coupling and dynamics of hybridized collective nematic-phonon excitations

### v1

@@ I Introduction
We find that the quasistatic picture above is qualitatively modified when considering the full dynamics of the coupled electronic-nematic-lattice degrees of freedom. Our main result is that the usual soft nematic collective mode illustrated in Fig. 1 – which in the absence of a lattice becomes underdamped as the QCP is approached [29] – remains damped even at the QCP once the lattice is included. Importantly, this mode remains the sharpest along the diagonal directions of the system, which is where the nematic form factor vanishes and the phonon velocity softens. Moreover, a new hybrid collective mode emerges which becomes progressively more coherent as the QCP is approached. This mode is a mixture of the original gapless nematic and phononic modes and, as a consequence of Adler’s theorem, it remains massless regardless of the distance to the QCP.
@@ III.2 Structural transitions and transverse phonons
We note that, since the matrix of elastic constants is $6\times 6$, the eigenvectors of the $3\times 3$ dynamical matrix do not generally produce the combinations of strains that are eigenvectors of the elastic constant matrix [31]. As $c_{11}-c_{12}\rightarrow 0$, the velocity of the modes propagating along $[110]$ and $[1\bar{1}0]$ vanishes. Importantly, these are transverse modes with polarization perpendicular to the direction of propagation. Hence, this tetragonal-to-orthorhombic structural transition is associated with a vanishing velocity of transverse phonons propagating along a diagonal direction.
@@ IV.3 Dispersion relations
Figure 9: Phonon dispersion at the QCP, $x=0$. Real (a) and imaginary (b) parts of the phonon dispersion for $\lambda_{\rm ph}/v_{s}=0.1$ plotted as a function of $q_{x}$ and $q_{y}$. (c)–(d) shows the real and imaginary parts of the dispersion as functions of $\theta_{\mathbf{q}}$ for fixed $q_{0}$ as $\lambda_{\rm ph}/v_{s}$ is varied. Here, the real part is normalized to the $\theta_{\mathbf{q}}=0$ value of the dispersion, $\tilde{v}_{s}q$, where $\tilde{v}_{s}=\sqrt{v_{s}^{2}+\lambda_{\rm ph}^{2}}$, while the imaginary part is normalized such that its minimum happens at $-1$. In (e)–(f), the real and imaginary parts of the nematic dispersion are shown as functions of $q$ for fixed $\theta_{\mathbf{q}}=\tfrac{\pi}{4}$. Note that the imaginary part for the $\lambda_{\rm ph}/v_{s}=0$ case vanishes identically. Crucially, the phonon mode exhibits the critical behavior seen in the nematic mode in the uncoupled case: For finite values of $\lambda_{\rm ph}$, the real part is quadratic in $q$ while the imaginary part is cubic in $q$. (g)–(h) shows the ratios between the real and the imaginary parts as a function of (g) $\theta_{\mathbf{q}}$ and (h) $q$.
@@ V Conclusions
In this work, we presented a derivation of the nemato-elastic coupling from a microscopic model, for which the coupling is mediated by the electronic degrees of freedom. In particular, the nematic fluctuations couple to the lattice vibrations indirectly via the coupling between electrons and transverse acoustic phonons. Since the latter vanishes in the simplest electron-phonon process due to the phonon’s transverse nature, we invoke the mechanism proposed in Ref.  [40, 41], which involves the impact of phonons on the impurity potential, Eq. (28). This formalism enables us to assess the hybridized nematic-phonon modes as a putative nematic QCP is approached. Due to the coupling to the gapless particle-hole excitations of the metal, an overdoped nematic mode that becomes underdoped at the QCP along certain directions emerges, governing the critical properties of the nematic instability in the absence of the lattice degrees of freedom. These nematic-plasmonic modes hybridize with the massless lattice modes – the transverse phonons – giving rise to two hybrid modes.
These hybrid modes exhibit a mixture of the dynamics characteristic of nematic fluctuations and of transverse phonons. Our systematic investigation of the on-shell dispersions of these hybrid modes as functions of both $q$ and $\theta_{\mathbf{q}}$ highlights distinct momentum-space regions of coherent and incoherent modes. Importantly, only one of the two hybrid modes exhibit the characteristic signatures of a critical mode, which is a consequence of Adler’s theorem, here manifesting in the massless modes.
@@ Appendix C Analytical approximation for the on-shell modes
To gain some analytical insight into the impact of $\lambda_{\rm ph}$ on the collective hybridized nematic-phonon modes, it is convenient to use the eigenvalues of Eq. (38) (in the absence of the $-\omega^{2}\mathds{1}$ part which will be reinstated later) to determine the poles. These modes are
@@
where, for convenience, $\widetilde{\chi}^{-1}_{\rm dres}(\mathbf{q},\omega)$ and $\widetilde{\mathcal{D}}^{-1}_{\rm dres}(\mathbf{q},\omega)$ refer to the dressed propagators without the $-\omega^{2}$-term:
@@
where, without loss of generality, we changed from $E_{\pm}$ to $E_{\rm nem}$ and $E_{\rm ph}$ to emphasize their relation to the unperturbed propagators. Here, $E_{\rm nem}$ originates from $E_{+}$ if $\widetilde{\chi}_{\rm dres}^{-1}(\mathbf{q},\omega)>\widetilde{\mathcal{D}}_{\rm dres}^{-1}(\mathbf{q},\omega)$ and from $E_{-}$ if $\widetilde{\chi}_{\rm dres}^{-1}(\mathbf{q},\omega)<\widetilde{\mathcal{D}}_{\rm dres}^{-1}(\mathbf{q},\omega)$ and vice versa for $E_{\rm ph}$. Focusing on the direction $\theta_{\mathbf{q}}=\tfrac{\pi}{4}$, we let $r\rightarrow x+\lambda_{\rm nem}^{2}\left(1+\frac{\lambda_{\rm ph}^{2}}{v_{s}^{2}-\lambda_{\rm ph}^{2}}\right)$ and expand Eqs. (56) and (57) to third order in $\tfrac{\omega}{v_{F}q}$ and reinstate the diagonal $-\omega^{2}$-term resulting in two equations for the dispersion of the nematic and phononic mode, respectively,

### v2

@@ I Introduction
We find that the quasistatic picture above is qualitatively modified when considering the full dynamics of the coupled electronic-nematic-lattice degrees of freedom. Our main result is that the usual soft nematic collective mode illustrated in Fig. 1 – which in the absence of a lattice becomes underdamped as the QCP is approached [29] – remains damped even at the QCP once the lattice is included. Importantly, this mode remains the sharpest along the diagonal directions of the system, which is where the nematic form factor vanishes and the phonon velocity softens. Moreover, a new hybrid collective mode emerges which becomes progressively more coherent as the QCP is approached. This mode is a mixture of the original gapless nematic and phononic modes and, as a consequence of Adler’s theorem, it remains massless regardless of the distance to the QCP. This behavior is consistent with recent observations of a lack of divergent effective masses near the nematic critical point in FeSe1-xSx [43, 44].
@@ III.2 Structural transitions and transverse phonons
We note that, since the matrix of elastic constants is $6\times 6$, the eigenvectors of the $3\times 3$ dynamical matrix do not generally produce the combinations of strains that are eigenvectors of the elastic constant matrix [31]. As $c_{11}-c_{12}\rightarrow 0$, the velocity of the modes propagating along $[110]$ and $1\bar{1}0]$ vanishes. Importantly, these are transverse modes with polarization perpendicular to the direction of propagation. Hence, this tetragonal-to-orthorhombic structural transition is associated with a vanishing velocity of transverse phonons propagating along a diagonal direction.
@@ IV.3 Dispersion relations
Figure 9: Phonon dispersion at the QCP, $x=0$. Real (a) and imaginary (b) parts of the phonon dispersion for $\lambda_{\rm ph}/v_{s}=0.1$ plotted as a function of $q_{x}$ and $q_{y}$. (c)–(d) shows the real and imaginary parts of the dispersion as functions of $\theta_{\mathbf{q}}$ for fixed $q_{0}$ as $\lambda_{\rm ph}/v_{s}$ is varied. Here, the real part is normalized to the $\theta_{\mathbf{q}}=0$ value of the dispersion, $\tilde{v}_{s}q$, where $\tilde{v}_{s}=\sqrt{v_{s}^{2}+\lambda_{\rm ph}^{2}}$, while the imaginary part is normalized such that its minimum happens at 1-1. In (e)–(f), the real and imaginary parts of the nematic dispersion are shown as functions of $q$ for fixed $\theta_{\mathbf{q}}=\tfrac{\pi}{4}$. Note that the imaginary part for the $\lambda_{\rm ph}/v_{s}=0$ case vanishes identically. Crucially, the phonon mode exhibits the critical behavior seen in the nematic mode in the uncoupled case: For finite values of $\lambda_{\rm ph}$, the real part is quadratic in $q$ while the imaginary part is cubic in $q$. (g)–(h) shows the ratios between the real and the imaginary parts as a function of (g) $\theta_{\mathbf{q}}$ and (h) $q$.
@@ V Conclusions
In this work, we presented a derivation of the nemato-elastic coupling from a microscopic model, for which the coupling is mediated by the electronic degrees of freedom. In particular, the nematic fluctuations couple to the lattice vibrations indirectly via the coupling between electrons and transverse acoustic phonons. Since the latter vanishes in the simplest electron-phonon process due to the phonon’s transverse nature, we invoke the mechanism proposed in Ref. [40, 41], which involves the impact of phonons on the impurity potential, Eq. (28). This formalism enables us to assess the hybridized nematic-phonon modes as a putative nematic QCP is approached. Due to the coupling to the gapless particle-hole excitations of the metal, an overdamped nematic mode that becomes underdamped at the QCP along certain directions emerges, governing the critical properties of the nematic instability in the absence of the lattice degrees of freedom. These nematic-plasmonic modes hybridize with the massless lattice modes – the transverse phonons – giving rise to two hybrid modes.
These hybrid modes exhibit a mixture of the dynamics characteristic of nematic fluctuations and of transverse phonons. Our systematic investigation of the on-shell dispersions of these hybrid modes as functions of both $q$ and $\theta_{\mathbf{q}}$ highlights distinct momentum-space regions of coherent and incoherent modes. Importantly, only one of the two hybrid modes exhibit the characteristic signatures of a critical mode, which is a consequence of Adler’s theorem, here manifesting in the massless modes. These results are consistent with the observed lack of divergent effective masses in S-doped FeSe near the nematic critical point [43, 44].
@@ Appendix C Analytical approximation for the on-shell modes
To gain some analytical insight into the impact of $\lambda_{\rm ph}$ on the collective hybridized nematic-phonon modes, it is convenient to use the eigenvalues of Eq. (38) (in the absence of the $\omega^{2}\mathds{1}$ part which will be reinstated later) to determine the poles. These modes are
@@
where, for convenience, $\widetilde{\chi}^{-1}_{\rm dres}(\mathbf{q},\omega)$ and $\widetilde{\mathcal{D}}^{-1}_{\rm dres}(\mathbf{q},\omega)$ refer to the dressed propagators without the $\omega^{2}-$term:
@@
where, without loss of generality, we changed from $E_{\pm}$ to $E_{\rm nem}$ and $E_{\rm ph}$ to emphasize their relation to the unperturbed propagators. Here, $E_{\rm nem}$ originates from $E_{+}$ if $\widetilde{\chi}_{\rm dres}^{-1}(\mathbf{q},\omega)>\widetilde{\mathcal{D}}_{\rm dres}^{-1}(\mathbf{q},\omega)$ and from $E_{-}$ if $\widetilde{\chi}_{\rm dres}^{-1}(\mathbf{q},\omega)<\widetilde{\mathcal{D}}_{\rm dres}^{-1}(\mathbf{q},\omega)$ and vice versa for $E_{\rm ph}$. Focusing on the direction $\theta_{\mathbf{q}}=\tfrac{\pi}{4}$, we let $r\rightarrow x+\lambda_{\rm nem}^{2}\left(1+\frac{\lambda_{\rm ph}^{2}}{v_{s}^{2}-\lambda_{\rm ph}^{2}}\right)$ and expand Eqs. (56) and (57) to third order in $\tfrac{\omega}{v_{F}q}$ and reinstate the diagonal $\omega^{2}-$term resulting in two equations for the dispersion of the nematic and phononic mode, respectively,